# Nonuniqueness for high-dimensional ideal MHD equations via differential inclusion

### v1

@@ 1. Introduction and main results
For every solution $(u,b,p)$ obtained in Theorem 1.2, there exist $2\leq m_{\ast}<\infty$ and a sequence of solutions $\{(u_{k},b_{k},p_{k},f_{k},g_{k})\}\subset C_{0}^{\infty}$ to the ideal MHD in (1.1) such that $\|(u_{k},b_{k},p_{k})\|_{L^{\infty}}$ remains uniformly bounded, and for any $2\leq m\leq m_{\ast}$,
@@
The rest of this article is structured as follows. In Section 2, we recast the ideal MHD equations in the form of a differential inclusion and construct highly oscillatory localized plane-wave solutions to the linearized relaxation system in every direction inside the wave cone. Section 3 begins by convexifying the nonlinear pointwise constraint in the differential inclusion framework. Then for any relaxed solution lying in the interior of this convex set, we add the highly oscillatory building blocks constructed in Section 2 in order to derive an energy growth inequality, which is the crucial step in iteratively producing a sequence of smooth relaxed solutions approximating a weak solution of the ideal MHD equations. Finally, we complete the proofs of the main results in Theorems 1.2 and 1.4 by means of the Baire category and convex integration methods in Section 4.
@@ 2. Relaxation of ideal MHD and localized plane wave
$(U,V)^{t}=\begin{pmatrix}U\\ V\end{pmatrix},\quad A^{t}(U,V)^{t}A=(A^{t}UA,A^{t}VA)^{t}=\begin{pmatrix}A^{t}UA\\ A^{t}VA\end{pmatrix}$.
@@
For each $A\in\mathscr{G}$ and any divergence-free matrix field $W\in\mathscr{M}$, the matrix field $T(y):=A^{t}\cdot W(A^{-t}y)\cdot A$ remains divergence-free.
@@
$\mathscr{H}=\big{\{}$ $E=(E_{ij}^{kl})\in(C^{\infty}(\mathbb{R}^{n+1}))^{(n+1)^{4}}|\,E_{ij}^{kl}=-E_{ji}^{kl},\,E_{ij}^{kl}=-E_{ij}^{lk},$
$E_{(n+1)i}^{(n+1)j}=0,\,1\leq i,j\leq n+1\big{\}}.$
For any $E,F\in\mathscr{H}$, define the linear operator $\mathscr{L}$ as follows:
$\mathscr{L}(E,F)^{t}=$ $(\mathscr{L}(E),\mathscr{L}(F))^{t}=\mathscr{L}\begin{pmatrix}E\\ F\end{pmatrix}=\begin{pmatrix}\mathscr{L}(E)\\ \mathscr{L}(F)\end{pmatrix}$
$=$ $\frac{1}{2}\begin{pmatrix}\big{(}\sum_{k,l}\partial_{kl}(E_{kj}^{il}+E_{ki}^{jl})\big{)}_{i,j=1,...,n+1}\vskip 2.15277pt\\ \big{(}\sum_{k,l}\partial_{kl}(F_{kj}^{il}+F_{ki}^{jl})\big{)}_{i,j=1,...,n+1}\end{pmatrix}$.
@@
For any $E,F\in\mathscr{H}$, we have $W:=\mathscr{L}(E,F)^{t}\in\mathscr{M}$ and $\mathrm{div}_{y}W=0.$
@@
Part 1. Consider the case when $\overline{W}\in\mathscr{M}$ satisfies $\overline{W}e_{1}=0$ and $\overline{W}e_{n+1}\neq 0$. For simplicity, write $\overline{W}=(\overline{U},\overline{V})^{t}$ for some $\overline{U}\in\mathscr{K}_{1}$ and $\overline{V}\in\mathscr{K}_{2}$. Since $\overline{W}e_{1}=0$, it is valid to take
$\begin{cases}E_{1i}^{1j}=E_{i1}^{j1}=-E_{i1}^{1j}=-E_{1i}^{j1}=\overline{U}_{ij}\frac{\sin(Ny_{1})}{N^{2}},\\ F_{1i}^{1j}=F_{i1}^{j1}=-F_{i1}^{1j}=-F_{1i}^{j1}=\overline{V}_{ij}\frac{\sin(Ny_{1})}{N^{2}},\end{cases}$
and let the rest of the entries be zero. Then we have $E,F\in\mathscr{H}$. Choose a smooth cutoff function $\psi$ with the property that $\|\psi\|_{L^{\infty}}\leq 1$, $\psi=1$ on $B_{1/2}(0)$, and $\mathrm{supp}\,\psi\subset B_{1}(0)$. Denote $W=\mathscr{L}(\psi E,\psi F)^{t}.$ Then from Lemma 2.3, we obtain that $W\in\mathscr{M}$ and $\mathrm{div}_{y}W=0.$ By a direct computation, we have
@@
$\|W-\psi\mathscr{L}(E,F)^{t}\|_{L^{\infty}}\leq C\|\psi\|_{C^{2}}\|(E,F)^{t}\|_{C^{1}}\leq\frac{C\|\psi\|_{C^{2}}}{N}<\delta,$
which, together with the fact that $\mathrm{supp}(\psi\mathscr{L}(E,F)^{t})\subset B_{1}(0)$ and $\psi\mathscr{L}(E,F)^{t}\in\sigma$, shows that $W$ takes values in the $\delta$-neighborhood of $\sigma$.
Part 2. The general case is treated by reduction to the preceding situation. Observe that there exists $\xi\in\mathbb{R}^{n+1}\setminus{0}$ such that $\overline{W}\xi=0$. This, in combination with the fact that $\overline{W}e_{n+1}\neq 0$, reads that the vectors $\xi$ and $e_{n+1}$ span a two-dimensional subspace. Pick a basis $\{\eta_{i}\}_{i=1}^{n+1}$ of $\mathbb{R}^{n+1}$ satisfying that $\eta_{1}=\xi$ and $\eta_{n+1}=e_{n+1}$. Then one can find $A\in\mathscr{G}$ such that $Ae_{i}=\eta_{i}$, i=1,,n+1i=1,...,n+1. Let $\widehat{W}=A^{t}\overline{W}A$. Then we obtain
$\widehat{W}_{(n+1)(n+1)}=$ $e_{n+1}^{t}\widehat{W}e_{n+1}=(Ae_{n+1})^{t}\overline{W}Ae_{n+1}=e_{n+1}^{t}\overline{W}e_{n+1}=0,$
@@ 3. Convex geometry and energy growth
Based on the binomial theorem and Young’s inequality, the proof of Lemma 3.5 was previously given in Lemma 2.1 of [28] in the case when $m$ is an integer. In fact, the inequality in (3.2) also remains true for non-integer values of $m$. This fact can be obtained from the convexity of the function $t^{m}$. To be specific, by letting $\theta=1-(1+\varepsilon)^{-1/(m-1)}$, it follows from a direct calculation that for any $0<\varepsilon<1$,
@@ 4.1. Baire category method
At this stage, we are ready to give the proofs of Theorems 1.2 and 1.4.
Proofs of Theorems 1.2 and 1.4.
Step 1. This step is to prove Theorem 1.2. In light of Lemmas 4.1 and 4.2, it suffices to prove that for any fixed $m\geq 2$, if $(u,b,M,Q,q)$ is a continuity point of $I_{m}$, then we have $|u(x,t)|=|b(x,t)|=1$ for almost all $(x,t)\in\Omega$. Since $I_{m}$ is continuous at the point $(u,b,M,Q,q)$, we can choose a weak*-convergent sequence $\{(u_{k},b_{k},M_{k},Q_{k},q_{k})\}\subset X$ approaching $(u,b,M,Q,q)$. With the aid of Lemma 3.7 and a standard diagonalization, we can extract another sequence $(\tilde{u}_{k},\tilde{b}_{k},\widetilde{M}_{k},\widetilde{Q}_{k},\tilde{q}_{k})$ converging weakly* to the same limit and satisfying
@@
Step 2. It remains to give the proof of Theorem 1.4. For every solution $(u,b,p)$ given by Step 1, there exists some $2\leq m_{\ast}<\infty$ such that for any $2\leq m\leq m_{\ast}$, $(u,b,M,Q,p)\in X$ is a continuity point of $I_{m}$, where
$M=u\otimes u-b\otimes b,\quad Q=b\otimes u-u\otimes b.$ (4.2)
Then one can find a sequence $\{(u_{k},b_{k},M_{k},Q_{k},p_{k})\}\subset X_{0}$ such that
$(u_{k},b_{k},M_{k},Q_{k},p_{k})\longrightarrow(u,b,M,Q,p)\quad\text{in $L^{m}$}.$ (4.3)
Take

### v2

# arXiv:2509.06866v2[math.AP] 01 Mar 2026

@@ 1. Introduction and main results
For every solution $(u,b,p)$ obtained in Theorem 1.2, there exist a sequence of solutions $\{(u_{k},b_{k},p_{k},f_{k},g_{k})\}\subset C_{0}^{\infty}$ to the ideal MHD in (1.1) such that $\|(u_{k},b_{k},p_{k})\|_{L^{\infty}}$ remains uniformly bounded, and for any $m\geq 2$,
@@
The rest of this article is structured as follows. In Section 2, we recast the ideal MHD equations in the form of a differential inclusion and construct highly oscillatory localized plane-wave solutions to the linearized relaxation system in every direction inside the wave cone. Section 3 begins by convexifying the nonlinear pointwise constraint in the differential inclusion framework. Then for any relaxed solution lying in the interior of this convex set, we add the highly oscillatory building blocks constructed in Section 2 in order to derive an energy growth inequality, which is the crucial step in iteratively producing a sequence of smooth relaxed solutions approximating a weak solution of the ideal MHD equations. In Section 4, we complete the proof of Theorem 1.2 by presenting two distinct approaches: the Baire category method and the convex integration scheme. Theorem 1.4 then follows as a direct consequence.
@@ 2. Relaxation of ideal MHD and localized plane wave
$(U,V)^{t}=\begin{pmatrix}U\\ V\end{pmatrix},\quad A^{t}(U,V)^{t}A=\begin{pmatrix}A^{t}UA\\ A^{t}VA\end{pmatrix}$.
@@
For each $A\in\mathscr{G}$ and any divergence-free matrix field $W\in\mathscr{M}$, the matrix field $T(y):=A^{t}W(A^{-t}y)A$ remains divergence-free.
@@
$\mathscr{H}_{1}=\big\{$ $E=(E_{ij}^{kl})\in(C^{\infty}(\mathbb{R}^{n+1}))^{(n+1)^{4}}|\,E_{ij}^{kl}=-E_{ji}^{kl},\,E_{ij}^{kl}=-E_{ij}^{lk},$
$E_{(n+1)i}^{(n+1)j}=0,\,1\leq i,j\leq n+1\big\},$
$\mathscr{H}_{2}=\big\{$ $E=(E_{ij}^{kl})\in\mathscr{H}_{1}|\,E^{il}_{kj}=0,\,1\leq i,j\leq n,1\leq k,l\leq n+1\big\},$
$\mathscr{H}_{3}=\big\{$ $E=(E_{ij}^{kl})\in(C^{\infty}(\mathbb{R}^{n+1}))^{(n+1)^{4}}|\,E_{ij}^{kl}=-E_{ji}^{kl},\,E_{ij}^{kl}=-E_{ij}^{lk},$
$E_{kj}^{(n+1)l}=E_{k(n+1)}^{jl}=0,\,1\leq j,k,l\leq n+1\big\}.$
For any $E\in\mathscr{H}_{1},F\in\mathscr{H}_{2},G\in\mathscr{H}_{3}$, define the linear operator $\mathscr{L}$ as follows:
$\mathscr{L}(E,F,G)=$ $\frac{1}{2}\begin{pmatrix}\big(\sum_{k,l}\partial_{kl}(E_{kj}^{il}+E_{ki}^{jl})\big)_{i,j=1,...,n+1}\vskip 2.15277pt\\ \big(\sum_{k,l}\partial_{kl}(F_{kj}^{il}+F_{ki}^{jl}+G_{kj}^{il}-G_{ki}^{jl})\big)_{i,j=1,...,n+1}\end{pmatrix}$.
@@
For any $E\in\mathscr{H}_{1},F\in\mathscr{H}_{2},G\in\mathscr{H}_{3}$, we have $W:=\mathscr{L}(E,F,G)\in\mathscr{M}$ and $\mathrm{div}W=0.$
@@
Part 1. Consider the case when $\overline{W}\in\mathscr{M}$ satisfies $\overline{W}e_{1}=0$ and $\overline{W}e_{n+1}\neq 0$. For simplicity, write $\overline{W}=(\overline{U},\overline{V})^{t}$ for some $\overline{U}\in\mathscr{K}_{1}$ and $\overline{V}\in\mathscr{K}_{2}$. Decompose $\overline{V}$ into two parts as follows:
$\overline{V}=\overline{V}_{\mathrm{sym}}+\overline{V}_{\mathrm{skew}},$
where
$\overline{V}_{\mathrm{sym}}\in\left\{\begin{pmatrix}0&b\\ b&0\end{pmatrix}\bigg|\,b\in\mathbb{R}^{n}\right\},\quad\overline{V}_{\mathrm{skew}}\in\left\{\begin{pmatrix}Q&0\\ 0&0\end{pmatrix}\bigg|\,Q\in\mathscr{S}_{-}^{n}\right\}.$
Since $\overline{W}e_{1}=0$, it is valid to take
$\begin{cases}E_{1i}^{1j}=E_{i1}^{j1}=-E_{i1}^{1j}=-E_{1i}^{j1}=\overline{U}_{ij}\frac{\sin(Ny_{1})}{N^{2}},\\ F_{1i}^{1j}=F_{i1}^{j1}=-F_{i1}^{1j}=-F_{1i}^{j1}=(\overline{V}_{\mathrm{sym}})_{ij}\frac{\sin(Ny_{1})}{N^{2}},\\ G_{1i}^{1j}=G_{i1}^{j1}=-G_{i1}^{1j}=-G_{1i}^{j1}=(\overline{V}_{\mathrm{skew}})_{ij}\frac{\sin(Ny_{1})}{N^{2}},\end{cases}$
and let the rest of the entries be zero. Then we have $E\in\mathscr{H}_{1},F\in\mathscr{H}_{2},G\in\mathscr{H}_{3}$. Choose a smooth cutoff function $\psi$ with the property that $\|\psi\|_{L^{\infty}}\leq 1$, $\psi=1$ on $B_{1/2}(0)$, and $\mathrm{supp}\,\psi\subset B_{1}(0)$. Denote $W=\mathscr{L}(\psi E,\psi F,\psi G).$ Then from Lemma 2.3, we obtain that $W\in\mathscr{M}$ and $\mathrm{div}W=0.$ By a direct computation, we have
@@
$\|W-\psi\mathscr{L}(E,F,G)\|_{L^{\infty}}\leq C\|\psi\|_{C^{2}}\|(E,F,G)\|_{C^{1}}\leq\frac{C\|\psi\|_{C^{2}}}{N}<\delta,$
which, together with the fact that $\mathrm{supp}(\psi\mathscr{L}(E,F,G))\subset B_{1}(0)$ and $\psi\mathscr{L}(E,F,G)\in\sigma$, shows that $W$ takes values in the $\delta$-neighborhood of $\sigma$.
Part 2. The general case is treated by reduction to the preceding situation. Observe that there exists $\xi\in\mathbb{R}^{n+1}\setminus\{0\}$ such that $\overline{W}\xi=0$. This, in combination with the fact that $\overline{W}e_{n+1}\neq 0$, reads that the vectors $\xi$ and $e_{n+1}$ span a two-dimensional subspace. Pick a basis $\{\eta_{i}\}_{i=1}^{n+1}$ of $\mathbb{R}^{n+1}$ satisfying that $\eta_{1}=\xi$ and $\eta_{n+1}=e_{n+1}$. Then one can find $A\in\mathscr{G}$ such that $Ae_{i}=\eta_{i}$, i=1,,n+1i=1,...,n+1. Let $\widehat{W}=A^{t}\overline{W}A$. Then we obtain
$e_{n+1}^{t}\widehat{W}e_{n+1}=(Ae_{n+1})^{t}\overline{W}Ae_{n+1}=e_{n+1}^{t}\overline{W}e_{n+1}=0,$
@@ 3. Convex geometry and energy growth
Based on the binomial theorem and Young’s inequality, the proof of Lemma 3.5 was previously given in Lemma 2.1 of [28] in the case when $m$ is an integer. In fact, the inequality in (3.2) also remains true for non-integer values of $m$. This fact can be obtained from the convexity of the power function. To be specific, by letting $\theta=1-(1+\varepsilon)^{-1/(m-1)}$, it follows from a direct calculation that for any $0<\varepsilon<1$,
@@ 4.1. Baire category method
A consequence of Lemma 4.2, Hölder’s inequality, the boundedness and compact support of functions in $(X,d^{\ast}_{\infty})$ gives the following result.
Corollary 4.3.
For any $m>2$, the set of continuity points of $I_{m}$ coincides with that of $I_{2}$.
We are now ready to give the proof of Theorem 1.2.
Proof of Theorem 1.2.
Combining Lemmas 4.1 and 4.2 with Corollary 4.3, it suffices to prove that for any fixed $m\geq 2$, if $(u,b,M,Q,q)$ is a continuity point of $I_{m}$, then we have $|u(x,t)|=|b(x,t)|=1$ for almost all $(x,t)\in\Omega$. Since $I_{m}$ is continuous at the point $(u,b,M,Q,q)$, we can choose a weak*-convergent sequence $\{(u_{k},b_{k},M_{k},Q_{k},q_{k})\}\subset X_{0}$ approaching $(u,b,M,Q,q)$. With the aid of Lemma 3.7 and a standard diagonalization, we can extract another sequence $(\tilde{u}_{k},\tilde{b}_{k},\widetilde{M}_{k},\widetilde{Q}_{k},\tilde{q}_{k})$ converging weakly* to the same limit and satisfying
@@ 4.2. Convex integration scheme
It remains to prove Theorem 1.4.
Proof of Theorem 1.4.
For every solution $(u,b,p)$ given by Theorem 1.2, it follows from the above proofs that there exists a sequence $\{(u_{k},b_{k},M_{k},Q_{k},p_{k})\}\subset X_{0}$ such that for any $m\geq 2$,
$(u_{k},b_{k},M_{k},Q_{k},p_{k})\longrightarrow(u,b,M,Q,p)\quad\text{in $L^{m}$},$ (4.4)
where
$M=u\otimes u-b\otimes b,\quad Q=b\otimes u-u\otimes b.$ (4.5)
Take